### v1

@@ 2.9 Implementing the PV installation
One of the critical issues to enable the physical installation of the system is related to the changes in the static load of the building. In our case, the local regulation establishes that changes in the static load of the rooftop above 5% require a building permit from the municipality, which could have delayed our process a few months. Hence, our first step was to obtain an engineering evaluation of the maximum load of the PV modules and structures on the rooftop that could be allowed below that threshold. This process was very time-consuming and included (i) consulting the municipality archives to retrieve the documents including a detailed description of the rooftop structure, (ii) contracting a small invasive examination of the rooftop by an engineering consultancy company to verify the documents.
Figure 6: Scheme of the PV installations in two different rooftops.
The previous steps allowed us to obtain the maximum weight that could be allowed on the rooftop and the areas on which the PV modules should be placed. This was taken into consideration when selecting among the offers received from different contractors. The board of the energy community assigned the task of evaluating the offer to a small technical committee comprised of members of the energy community. In the evaluation, the criteria for reliability (e.g. selecting a proven configuration) was prioritized over the cost. A delta configuration including two rows of PV modules facing east and west respectively, and tilted 10º was selected (see Fig. 5). This was preferred over a classic south-oriented installation for the following reason. Despite a slightly lower annual generation, the east-west delta configuration reduces the wind loads and, consequently, the need for ballast and weight of the installation. Following the criteria to maximize system reliability over cost, a ballast system was chosen over a glued-structure system because the latter was considered less-proven. Selecting the placement of the inverters also took some discussion. The final agreement was to place them in the building’s technical room where they are protected from outdoor weather.
Since our building is in reality comprised of several towers, the final system includes two subsystems, each of them with its own inverter, see Fig. 6. We took this opportunity to select two types of PV modules (monofacial and bifacial monocrystalline silicone) and will use our installation to evaluate which of the two performs better under the Danish weather.
@@ 2.10 Operating the installation
Regarding the selling of electricity to the grid for hours in which the PV generation is higher than the local demand, we are currently evaluating several options. One difficulty is that currently, the electricity market in Denmark allows negative prices, which are typically observed in hours with high renewable generation, e.g. during sunny middays. If we were to sign a contract to sell the excess electricity to the grid, the current Danish regulation establishes that we should also pay when negative prices happen. This poses a risk in our strategy since we could potentially lose money by exporting electricity to the grid and it might be safer to just export the electricity to the grid for free without signing a contract. Moreover, most of the companies which could buy our excess electricity require paying certain grid tariffs and handling charges, which reduces the economic benefit of exported electricity. The initial business plan assumed that 85% of the energy would be locally consumed by the university and the remaining would be exported for free. Hence, whether to export surplus electricity for free and who will we sell to would also impact our business model and recovery time.
@@ 3 Conclusions and policy recommendations
At the national energy regulation in Denmark, policy recommendations include (i) establishing legally the possibility for universities to become energy communities, (ii) removing the electricity tax (‘elafgift’) on the locally-produced electricity under the scheme of ´self-consumption via third party’ when the owner of the installation or the consumer of the electricity is an energy community (Table 2.1), (iii) relieving the energy communities of the need to pay negative electricity prices when they export electricity to the grid. On top of all the social and environmental benefits brought by energy communities, the second proposal is an extension of the electricity tax exemption currently only valid for normal consumers under self-consumption regulation. The third suggestion would improve the economic viability of energy communities by reducing their associated economic risk at a very low cost for society.
At the municipal level, energy communities can be fostered by (i) creating the public of energy-community-enabler public employees as described in Section 2.11, (ii) guaranteeing easy access to building data for a fast evaluation of static load changing for new rooftop PV installation, (iii) creating an easy-to-use platform to crowdsource the fundings for energy communities.
Energy communities can be the strategy to enable collective and citizen-driven energy actions to support the clean energy transition, but for this to be true it is needed to simplify the administrative work, set up reasonable economic frameworks and collect the know-how and experience through online repositories and energy-community-enabler public employees.

### v2

@@ 2.9 Implementing the PV installation
One of the critical issues to enable the physical installation of the system is related to the changes in the static load of the building. In our case, the local regulation establishes that changes in the static load of the rooftop above 5% require a building permit from the municipality, which could have delayed our process for a few months. Hence, our first step was to obtain an engineering evaluation of the maximum load of the PV modules and structures on the rooftop that could be allowed below that threshold. This process was very time-consuming and included (i) consulting the municipality archives to retrieve the documents including a detailed description of the rooftop structure, (ii) contracting a small invasive examination of the rooftop by an engineering consultancy company to verify the documents.
Figure 6: Scheme of the PV installations on two different rooftops.
The previous steps allowed us to obtain the maximum weight that could be allowed on the rooftop and the areas on which the PV modules should be placed. This was taken into consideration when selecting among the offers received from different contractors. The board of the energy community assigned the task of evaluating the offer to a small technical committee comprised of members of the energy community. In the evaluation, the criteria for reliability (e.g. selecting a proven configuration) was prioritized over the cost. A delta configuration including two rows of PV modules facing east and west respectively, and tilted 10º was selected (see Fig. 5). This was preferred over a classic south-oriented installation for the following reasons. Despite a slightly lower annual generation, the east-west delta configuration reduces the wind loads and, consequently, the need for ballast and weight of the installation. Following the criteria to maximize system reliability over cost, a ballast system was chosen over a glued-structure system because the latter was considered less-proven. Selecting the placement of the inverters also took some discussion. The final agreement was to place them in the building’s technical room, where they are protected from outdoor weather.
Since our building is in reality comprised of several towers, the final system includes two subsystems, each of them with its own inverter, see Fig. 6. We took this opportunity to select two types of PV modules (comprising monofacial and bifacial monocrystalline silicon solar cells) and will use our installation to evaluate which of the two performs better under the Danish weather. The installation was completed and the system started producing electricity in October 2024.
@@ 2.10 Operating the installation
Regarding the selling of electricity to the grid for hours in which the PV generation is higher than the local demand, we are currently evaluating several options. One difficulty is that currently, the electricity market in Denmark allows negative prices, which are typically observed in hours with high renewable generation, e.g. during sunny middays. If we were to sign a contract to sell the excess electricity to the grid, the current Danish regulation establishes that we should also pay when negative prices happen. This poses a risk in our strategy since we could potentially lose money by exporting electricity to the grid and it might be safer to just export the electricity to the grid for free without signing a contract. Moreover, most of the companies which could buy our excess electricity require paying certain grid tariffs and handling charges, which reduces the economic benefit of exported electricity. The initial business plan assumed that 85% of the energy would be locally consumed by the university and the remaining would be exported for free. Hence, whether to export surplus electricity for free and to whom we will sell would also impact our business model and recovery time.
@@ 2.11 Replicating our experience
3 Putting our challenges into perspective. Similarities with existing literature
Many researchers have previously attempted to summarize the challenges and barriers associated with implementing energy communities. These challenges span across regulatory, financial, technical, and organizational dimensions. In this section, we summarize previous literature and reflect on where it aligns or contradicts our experience.
Across several studies in Europe, regulatory and policy barriers are frequently cited as the top obstacle to energy community implementation [11, 12, 13, 14, 15, 16]. In some cases, the issue is the lack of clear regulatory and legal framework supporting the creation and operation of energy communities [11, 12, 13, 14, 15, 16]. Dioba et al. [15], however, emphasize that the complexity of regulation, more so than legal limitations, is the most significant barrier across European Union countries. In our case, it took us some effort and required help from an external lawyer to set up the legal framework for the energy community confirming previous literature.
Financial and economic constraints also play a critical role. Factors such as high capital cost and limited access to financing mechanisms are commonly identified in the literature [11, 12, 14, 15, 16]. In particular, several studies highlight the difficulty in securing bank loans for community energy projects[11, 16]. Coupled with uncertain or slow returns on investment [13, 14] and market uncertainties [15], these financial and economic constraints can significantly limit the implementation of energy community projects. In our case, we limited the value of one share to 900 DKK (120 EUR) to facilitate funding and participation. The drawback is the higher administrative costs to manage a cooperative composed of many members with low investment per share. The market uncertainty identified in previous literature has been clearly a large challenge in our case, as for example, the risk of negative prices has prevented the cooperative from establishing a contract to export electricity to the grid.
A few studies identify the reliance on volunteers as a challenge for energy communities [11, 12, 17, 18]. This can lead to both organizational and technical challenges. From an administrative perspective, volunteers typically have limited time to dedicate to projects, resulting in constrained human resources [12, 17]. Technically, this translates to a lack of expertise and competencies [17]. As a result, communities may need to hire external experts, which in turn raises project costs and adds to the financial challenge [11]. Here, we have benefited from European funding, via the AURORA project, to hire a community coordinator, as well as benefit from large voluntary work from members. The hiring of a community coordinator was key to kickstart the project.
Additionally, both Aoidh et al. [16] and Dioba et al. [15] , point to low trust and mistrust in the community model as further challenges. A few sources cite that lack of information, whether technical, legal, or financial, as a hurdle [19, 20]. However, De Franco et al technical, legal, or financial, as a hurdle [19] argue that it is not necessarily the lack of comprehendible information but rather how information is interchanged and communicated that challenges community engagement. In our experience, this has not been a barrier, but the community members showed a high level of trust in the board of the energy community. The plans and actions taken were communicated through a mailing group and general assemblies are held annually.
@@ 4 Conclusions and policy recommendations
At the national energy regulation in Denmark, policy recommendations include (i) establishing legally the possibility for universities to become energy communities, (ii) removing the electricity tax (‘elafgift’) on the locally-produced electricity under the scheme of self-consumption via third party when the owner of the installation or the consumer of the electricity is an energy community (Table 1), and (iii) relieving the energy communities of the need to pay negative electricity prices when they export electricity to the grid. On top of all the social and environmental benefits brought by energy communities, the second proposal is an extension of the electricity tax exemption currently only valid for normal consumers under self-consumption regulation. The third suggestion would improve the economic viability of energy communities by reducing their associated economic risk at a very low cost for society.
At the municipal level, energy communities can be fostered by (i) creating the figure of energy-community-enabler public employees as described in Section 2.11, (ii) guaranteeing easy access to building data for a fast evaluation of static load for new rooftop PV installation, (iii) creating an easy-to-use platform to crowdsource the fundings for energy communities.
Energy communities can be the strategy to enable collective and citizen-driven energy actions to support the clean energy transition, but for this to be true, it is needed to simplify the administrative work, set up reasonable economic frameworks and collect the know-how and experience through online repositories and energy-community-enabler public employees.